# SAM: A Stability-Aware Cache Manager for Multi-Tenant Embedded Databases

### v1

@@ Abstract.
The co-location of multiple database instances on resource constrained edge nodes creates significant cache contention, where traditional schemes are inefficient and unstable under dynamic workloads. To address this, we present SAM, an autonomic cache manager powered by our novel AURA algorithm. AURA makes stability a first-class design principle by resolving the exploitation-exploration dilemma: it achieves this by synthesizing two orthogonal factors, which we introduce as: the $\mathcal{H}$-factor, representing a database’s proven, historically stable efficiency (exploitation), and the $\mathcal{V}$-factor, representing its empirically estimated marginal gain for future improvements (exploration). This dual-factor model, governed by an adaptive weight, enables SAM to achieve sustained high performance through strategic stability and robustness in volatile conditions.
Extensive experiments against 14 diverse baselines demonstrate SAM’s superiority. It achieves top-tier throughput while being uniquely resilient to complex workload shifts and cache pollution attacks. Furthermore, its decision latency is highly scalable, remaining nearly constant as the system grows to 120 databases. Crucially, SAM achieves superior decision stability—maintaining consistent optimization directions despite noise, avoiding performance oscillations while ensuring predictable Quality of Service. These results prove that a principled, stability-aware design is essential for sustained high performance in real-world, large-scale systems.
@@ 1. Introduction
To effectively balance this exploitation-exploration trade-off, we propose SAM (a Stability-Aware Manager), an autonomic cache management system powered by our novel AURA (Autonomic Utility-balancing Resource Allocator) algorithm. As illustrated in Figure 1, SAM’s design directly maps onto the challenges. First, its two-pool allocation model institutionalizes the “reservation plus elasticity” paradigm, providing deterministic baseline resources via a fixed pool to directly address (C3) SLA Guarantees. Second, and most critically, the AURA algorithm provides the principled, dual-factor scoring model for the elastic pool. It balances a database’s historically stable efficiency ($\mathcal{H}$-factor) to prevent (C2) Cache Pollution, with its estimated marginal gain ($\mathcal{V}$-factor) to adapt to (C1) Workload Dynamism. Finally, AURA’s stability-enforcing actuation layer ensures (C4) Decision Stability. Together, these mechanisms enable SAM to achieve sustained high performance by combining decision stability with robustness against external disturbances.
@@
    We propose SAM, a novel cache management system for multi-tenant embedded databases. Its two-pool architecture is the first to systematically decouple baseline SLA guarantees from dynamic performance optimization.
@@
    We propose the AURA algorithm, whose novelty lies in a dual-factor model that balances a tenant’s historically stable efficiency ($\mathcal{H}$-factor) with its forward-looking marginal gain ($\mathcal{V}$-factor). This principled synthesis achieves decision stability that prevents self-inflicted performance oscillations and robustness against deceptive workloads, together enabling sustained high performance in volatile environments.
@@
    Through extensive evaluation against 14 diverse baselines, we demonstrate that SAM achieves top-tier throughput while maintaining unique robustness to adversarial workloads like cache pollution. Our scalability analysis shows near-constant decision latency as the system scales to 120 databases.
The remainder of this paper is organized as follows. Section 2 presents the system model and problem formulation. Section 3 details the SAM system’s architecture and design philosophy. Section 4 presents the AURA scheduling algorithm in detail. Section 5 provides a theoretical analysis of the algorithm’s stability properties and its computational complexity. Section 6 describes our experimental setup and evaluation. Section 7 reviews related work, and Section 8 concludes.
@@ 2.3. Conflicting Design Goals
Exploitation vs. Exploration. This trade-off is central to addressing both Challenge 1 (Workload Dynamism) and Challenge 2 (Cache Pollution). An algorithm that solely exploits the current state may achieve high current throughput but risks missing emerging hotspots. Conversely, aggressive exploration by chasing naive metrics like miss count can be deceived by cache-unfriendly workloads. As detailed in Section 4, AURA resolves this by synthesizing an exploitation signal (the $\mathcal{H}$-factor) with an exploration signal (the $\mathcal{V}$-factor) via a meta-adaptive policy.
@@
Responsiveness vs. Decision Stability. This is the core trade-off for Challenge 4 (Decision Stability), especially in noisy edge environments. Overly reactive strategies are vulnerable to transient noise, causing performance oscillations that degrade sustained throughput. Conversely, excessive damping may miss genuine workload shifts. AURA’s stability-aware mechanisms, including its Saturation Confidence model and momentum-based allocator (Section 4), maintain decision consistency while preserving necessary adaptiveness.
@@ 3. System Overview
To address the constrained optimization problem defined in Section 2, we propose SAM, an autonomic cache management framework. The core of our stability-aware design philosophy is that principled, internal stability is the foundation for achieving both high performance and external robustness in volatile, multi-tenant environments. It moves beyond simple reactive heuristics by implementing our novel AURA algorithm, which achieves sustained high performance through a dual-factor model that balances historical efficiency with future potential.
@@ 3.1. Overall Architecture
The System Layer contains the core control logic. A centralized SAM Autonomic Coordinator, powered by our AURA algorithm, manages the database instances residing in the Managed Environment. This architecture implements a classic feedback control loop operating in a Sense-Decide-Act cycle:
Sense: The Coordinator gathers fine-grained Performance Metrics directly from each database instance, while also perceiving global System Info from the resource layer to maintain a comprehensive view of the system state. Decide: In the decision phase, the Coordinator executes the AURA algorithm. It synthesizes the perceived state to calculate an optimal allocation plan that balances historical performance with future potential, as will be detailed in Section 4. Act: The Coordinator dispatches direct Cache Resizing Commands to each target instance, which then adjusts its memory footprint by requesting resources from or releasing them to the Elastic Pool.
@@
Figure 1. The architecture of the SAM system, illustrating its two-pool resource model and Sense-Act control loop. A central SAM Autonomic Coordinator perceives metrics from the database instances (Sense) and issues direct resizing commands back to them (Act), managing a partitioned Fixed Pool and Elastic Pool.
@@ 3.2. The Two-Pool Cache Architecture
Fixed Pool ($C_{fixed}$). This pool embodies our architectural principle that deterministic performance baselines for critical instances should be guaranteed through static allocation, not pursued as a goal within the dynamic optimization loop. This guarantee is realized through a two-step, policy-driven allocation process. First, an administrator defines the total size of the fixed pool, $C_{fixed}$, based on overall system service-level objectives. Second, this pool is apportioned among instances ($d_{j}\in D_{p}$) proportionally to their static, pre-defined ,manually configured base_priority values. This static allocation, $a_{j}^{\text{fixed}}$, insulates critical instances from the resource volatility of the wider system. Critically, this separation of concerns allows the manager for the remaining Elastic Pool to focus single-mindedly on maximizing overall system utility, without being constrained by the safety requirements of individual tenants.
Elastic Pool ($C_{elastic}$). This pool contains all remaining cache resources ($C_{elastic}=C_{total}-C_{fixed}$) and is dedicated to maximizing the aggregate system utility, as defined by our objective function. The resources in this pool are entirely fluid and are dynamically reallocated by our core AURA algorithm, which will be introduced in the next section, in each tuning interval among all participating database instances. Finally, total allocation for any instance $d_{i}$ is the sum of its parts: $a_{i}^{\text{total}}(t)=a_{i}^{\text{fixed}}+a_{i}^{\text{elastic}}(t)$.
@@ 3.3. The SAM Coordinator and the AURA Control Policy
The design of our control policy is guided by two foundational principles. First, it is based on a dual-factor utility paradigm that resolves the exploitation-exploration trade-off by evaluating each instance along two orthogonal dimensions: its proven historical contribution (exploitation) and its marginal potential for future improvement (exploration). Second, it treats stability as a first-class citizen. Here, we distinguish between decision stability—maintaining consistent optimization directions despite noisy observations—and mere response damping that would slow adaptation. AURA employs control-theoretic mechanisms to achieve the former, dampening internal oscillations while ensuring robust adaptations against external disturbances.
@@
Our novel AURA algorithm, whose detailed specification is the subject of the next section, is the concrete implementation of this control policy. By synergistically synthesizing these principles, AURA provides the autonomic control policy that allows SAM to make stable, robust, and high-performance decisions, overcoming the limitations of simpler baselines.
@@
Figure 2. The two-level architecture of the AURA algorithm, showing: (a) the macro-level state machine that transitions between a fast-path Localized Optimization and an infrequent Global Scan to ensure scalability, and (b) the micro-level decision pipeline executed within the fast path.
4. The AURA Algorithm
@@
The design principles articulated in the previous section culminate in the AURA algorithm, the principled, dual-factor control core of our SAM framework. AURA is architected as a two-level, adaptive state management algorithm that implements our stability-aware control policy while maintaining computational efficiency suitable for large-scale deployments. Figure 2 provides a complete visual schematic of this architecture.
4.1. Architectural Principles of the AURA Algorithm
@@
As established in Section 2, the online, black-box nature of our problem makes finding a provably optimal solution intractable. Therefore, our approach is to design an autonomic control algorithm, AURA. AURA navigates the fundamental trade-off between exploitation and exploration to achieve the stability and robustness necessary for sustained high performance. It materializes this principle by synthesizing two orthogonal signals—a database’s proven, historical efficiency and its potential for future performance gains—allowing it to make sound decisions in complex, dynamic environments.
@@
At a macro-level, as shown in Figure 2, AURA employs an Adaptive Active Set (AAS) framework, which operates as a two-state machine. The algorithm spends the majority of its cycles in the computationally efficient State 1 (Localized Optimization), focusing only on a small subset of tenants ($k\ll K$) to make rapid, high-quality decisions. It transitions to the more comprehensive State 2 (Global Scan) only when a strategic opportunity for improvement is detected (i.e., the active set converges). This design decouples decision latency from the system scale, providing excellent scalability while maintaining decision quality.
@@ 4.2. Algorithm Specification
We now provide the complete specification of the AURA algorithm, translating the high-level architecture from the previous section into a concrete implementation. We begin by laying out the necessary data structures and state variables, followed by a detailed walkthrough of the main decision workflow and its innovative components as visually summarized in Figure 2.
@@
The AURA algorithm maintains state at two levels. The Global Coordinator State ($S_{g}$) includes: (i) Active Set State, which contains the current_active_set of tenants being optimized and cached candidate pools (top_candidates, bottom_candidates); (ii) Convergence State, such as a score_impro-
@@
The algorithm’s autonomic control policy employs an adaptive two-state architecture for computational efficiency. At the start of each cycle, a key decision is made (Algorithm 1, Line 1): should it enter the infrequent, comprehensive Global Scan state, or continue in the lightweight, default Localized Optimization state? This decision is governed by a convergence metric that determines when the optimization potential of the current active set has been exhausted.
If a global scan is triggered (e.g., on a cold start or after convergence), the system performs a highly efficient, linear-time pass using a two-way heap filter to gather fresh candidates (Line 1). After passing an equilibrium check, a knee-point detection algorithm robustly determines the optimal size for a new active set (Line 1).
However, the majority of cycles are spent in the fast-path Localized Optimization state. Here, the computationally intensive scoring and allocation adjustments are confined to the small, $O(k)$ active set via the OptimizeInActiveSet function (Line 1). This two-state architecture is the key to AURA’s ability to provide high-quality decisions at a near-constant low cost, regardless of system scale. The core components used within the OptimizeInActiveSet function—the dual-factor scoring model and its supporting mechanisms—are detailed below.
@@ 9    $k_{\text{demand}}\leftarrow\operatorname{FindKneePoint}(TopK)$
Algorithm 1 The AURA Dual-Factor Control Algorithm
@@
The Meta-Adaptive Policy ($\alpha_{t}$). To dynamically balance the $\mathcal{H}$-factor (exploitation) and $\mathcal{V}$-factor (exploration), AURA employs a meta-adaptive policy that computes the weight $\alpha_{t}$. This policy acts as a system-wide “confidence dial”, driven by the insight that the statistical variance of $\mathcal{V}$-factors across the active set serves as a robust proxy for environmental uncertainty. A normalized variance metric, $\kappa$, is calculated; a high $\kappa$ signifies a volatile state with contradictory signals, pushing the system to be more conservative by increasing the weight on the stable $\mathcal{H}$-factor. Conversely, a low $\kappa$ indicates a stable state with reliable signals, allowing the system to be more opportunistic by trusting the $\mathcal{V}$-factor. The final value is temporally smoothed to ensure the policy itself remains stable.
@@
To understand how the composite scoring function translates into behavior, we analyze its response to four key database archetypes. (i) The Saturated, High-Contribution Instance: This instance maintains a high score due to its large $\mathcal{H}$-factor (from high ops and hit rate). However, its $\mathcal{V}$-factor converges to zero as additional cache yields no further improvement ($\Delta HR\approx 0$), correctly preventing it from wastefully accumulating more resources. (ii) The High-Potential, Low-Resource Instance: An emerging hotspot is characterized by a high marginal gain ($\mathcal{V}$-factor), as small cache additions yield large hit rate improvements. Our normalization mechanism identifies this strong signal, boosting its score and allowing it to rapidly acquire the resources it needs. (iii) The Cache-Inefficient Workload: A “cache polluter” with high ops but persistently low hit rate receives a near-zero score. This is because its $\mathcal{H}$-factor is suppressed by the low hr_norm, and its $\mathcal{V}$-factor is zero, leading to the swift reclamation of its cache. (iv) The Quiescent Instance: An idle database has near-zero ops and no activity, causing both its $\mathcal{H}$-factor and $\mathcal{V}$-factor to be zero. Consequently, it receives no elastic cache, preserving resources for active instances.

### v2

@@ Abstract.
The co-location of multiple database instances on resource constrained edge nodes creates significant cache contention, where traditional schemes are inefficient and unstable under dynamic workloads. To address this, we present SAM(a Stability-Aware Manager), an autonomic cache manager that establishes decision stability as a first-class design principle. It achieves this through its core control policy, AURA(Autonomic Utility-balancing Resource Allocator), which resolves the classic exploitation-exploration dilemma by synthesizing two orthogonal factors: the $\mathcal{H}$-factor, representing proven historical efficiency (exploitation), and the $\mathcal{V}$-factor, for estimated marginal gain (exploration). Through this practical synthesis and adaptive control, SAM achieves sustained high performance with strategic stability and robustness in volatile conditions.
Extensive experiments against 14 diverse baselines demonstrate SAM’s superiority. It achieves top-tier throughput while being uniquely resilient to complex workload shifts and adversarial workloads like cache pollution. Furthermore, its decision latency is highly scalable, remaining nearly constant as the system grows to 120 databases. Crucially, SAM achieves superior decision stability—maintaining consistent optimization directions despite noise, avoiding performance oscillations while ensuring predictable Quality of Service. These results prove that a principled, stability-aware design is essential for sustained high performance in real-world, large-scale systems.
@@ 1. Introduction
To effectively balance this exploitation-exploration trade-off, we propose SAM (a Stability-Aware Manager), a closed-loop control system whose design directly addresses these challenges (Fig. 1). First, to ensure (C3) SLA Guarantees, SAM’s two-pool allocation model provides deterministic baseline resources via a fixed pool. Second, to tackle the dynamic challenges, SAM relies on its principled core control policy, AURA(Autonomic Utility-balancing Resource Allocator), to manage the elastic pool. Through a practical synthesis of two orthogonal factors, SAM prevents (C2) Cache Pollution by rewarding historically stable efficiency ($\mathcal{H}$-factor), while adapting to (C1) Workload Dynamism by estimating marginal gain ($\mathcal{V}$). Finally, to guarantee (C4) Decision Stability, SAM utilizes AURA’s stability-enforcing actuation layer. By unifying these design principles, SAM achieves sustained high performance, combining strategic stability with robustness against external disturbances.
@@
    We propose SAM, a cache management system built on a novel, stability-first design principle. Its two-pool architecture is the first to systematically decouple baseline SLA guarantees from dynamic performance optimization, architecturally separating safety from efficiency.
@@
    We introduce a principled engineering approach to stability-aware control, embodied in our policy, AURA. It navigates the complex exploitation-exploration trade-off by practically synthesizing a tenant’s historical efficiency ($\mathcal{H}$-factor) with its forward-looking marginal gain ($\mathcal{V}$-factor).
@@
    We provide extensive experimental validation of SAM’s superiority and practical viability against 14 diverse baselines. Our results on standard, adversarial, and TPC-C workloads demonstrate near-optimal throughput, unique resilience, and excellent scalability.
The remainder of this paper is organized as follows. Section 2 presents the system model and problem formulation. Section 3 details the SAM system’s architecture and design philosophy. Section 4 presents the AURA control policy in detail. Section 5 provides a theoretical analysis of the algorithm’s stability properties and its computational complexity. Section 6 describes our experimental setup and evaluation. Section 7 reviews related work, and Section 8 concludes.
@@ 2.3. Conflicting Design Goals
Exploitation vs. Exploration. This trade-off is central to addressing both Challenge 1 (Workload Dynamism) and Challenge 2 (Cache Pollution). An algorithm that solely exploits the current state may achieve high current throughput but risks missing emerging hotspots. Conversely, aggressive exploration by chasing naive metrics like miss count can be deceived by cache-unfriendly workloads. As detailed in Section 4, our system, SAM, resolves this through its core control policy, AURA, which practically synthesizes an exploitation signal (the $\mathcal{H}$-factor) with an exploration signal (the $\mathcal{V}$-factor).
@@
Responsiveness vs. Decision Stability. This is the core trade-off for Challenge 4 (Decision Stability), especially in noisy edge environments. Overly reactive strategies are vulnerable to transient noise, causing performance oscillations that degrade sustained throughput. Conversely, excessive damping may miss genuine workload shifts. Our system, SAM, is designed to maintain decision consistency while preserving necessary adaptiveness, utilizing the stability-aware mechanisms within its AURA policy, such as its Saturation Confidence model and momentum-based allocator (Section 4).
@@ 3. System Overview
To address the constrained optimization problem defined in Section 2, we propose SAM, an autonomic cache management framework. The core of our stability-aware design philosophy is that principled, internal stability is the foundation for achieving both high performance and external robustness in volatile, multi-tenant environments. SAM puts this philosophy into practice through its core control policy, AURA, which moves beyond simple reactive heuristics through a principled synthesis of historical efficiency ($\mathcal{H}$-factor) and future potential ($\mathcal{V}$-factor).
@@ 3.1. Overall Architecture
The System Layer contains the core control logic. A centralized SAM Autonomic Coordinator, which implements the AURA control policy, manages the database instances residing in the Managed Environment. This architecture implements a classic feedback control loop operating in a Sense-Decide-Act cycle:
Sense: The Coordinator gathers fine-grained Performance Metrics directly from each database instance, while also perceiving global System Info from the resource layer to maintain a comprehensive view of the system state. Decide: In the decision phase, the Coordinator executes its AURA control policy, synthesizing the perceived state to calculate an optimal allocation plan. This plan, detailed in Section 4, strategically balances historical performance with future potential. Act: The Coordinator dispatches direct Cache Resizing Commands to each target instance, which then adjusts its memory footprint by requesting resources from or releasing them to the Elastic Pool.
@@
Figure 1. Architecture of the SAM system, featuring a two-pool (Fixed/Elastic) resource model where a central Coordinator manages the elastic pool via a Sense-Decide-Act feedback loop.
@@ 3.2. The Two-Pool Cache Architecture
Fixed Pool ($C_{fixed}$). This pool embodies our architectural principle that deterministic performance baselines for critical instances should be guaranteed through static allocation, not pursued as a goal within the dynamic optimization loop. This guarantee is realized through a two-step, policy-driven allocation process. First, an administrator defines the total size of the fixed pool, $C_{fixed}$, based on overall system service-level objectives. Second, this pool is apportioned among instances ($d_{j}\in D_{p}$) proportionally to their static, pre-defined, manually configured base_priority values. The manually configured Fixed Pool is intentional, separating high-level policy set by administrators from the low-level, dynamic optimization handled by AURA. This static allocation, $a_{j}^{\text{fixed}}$, insulates critical instances from the resource volatility of the wider system. Critically, this same separation of concerns allows the manager for the remaining Elastic Pool to focus single-mindedly on maximizing overall system utility, without being constrained by the safety requirements of individual tenants.
Elastic Pool ($C_{elastic}$). This pool contains all remaining cache resources ($C_{elastic}=C_{total}-C_{fixed}$) and is dedicated to maximizing the aggregate system utility, as defined by our objective function. The resources in this pool are entirely fluid and are dynamically reallocated in each tuning interval by the SAM system, guided by its core AURA control policy (which will be introduced in the next section). Finally, total allocation for any instance $d_{i}$ is the sum of its parts: $a_{i}^{\text{total}}(t)=a_{i}^{\text{fixed}}+a_{i}^{\text{elastic}}(t)$.
@@ 3.3. The SAM Coordinator and the AURA Control Policy
The design of our control policy is guided by two foundational principles. First, it is based on a dual-factor utility paradigm that resolves the exploitation-exploration trade-off by evaluating each instance along two orthogonal dimensions: its proven historical contribution (exploitation) and its marginal potential for future improvement (exploration). Second, it treats stability as a first-class citizen. Here, we distinguish between decision stability—maintaining consistent optimization directions despite noisy observations—and mere response damping that would slow adaptation. Our policy employs control-theoretic mechanisms to achieve the former, dampening internal oscillations while ensuring robust adaptations against external disturbances.
@@
The concrete implementation of this control policy is our AURA engine, whose detailed specification is the subject of the next section. By effectively integrating these principles, this policy equips SAM with the ability to make stable, robust, and high-performance decisions, overcoming the limitations of simpler baselines.
@@
Figure 2. The two-level architecture of the AURA control policy, showing: (a) the macro-level state machine that transitions between a fast-path Localized Optimization and an infrequent Global Scan to ensure scalability, and (b) the micro-level decision pipeline executed within the fast path.
4. The AURA Control Policy
@@
The design principles articulated in the previous section are embodied in our core control policy, AURA, which serves as the control core of the SAM framework. This policy is realized through a two-level, adaptive state management architecture that maintains computational efficiency suitable for large-scale deployments. Figure 2 provides a visual schematic of this architecture.
4.1. Architectural Principles of the AURA
@@
As established in Section 2, the online, black-box nature of our problem makes finding a provably optimal solution intractable. Therefore, our approach is to design an autonomic control policy, AURA, to govern the SAM framework. This policy navigates the fundamental trade-off between exploitation and exploration to achieve the stability and robustness necessary for sustained high performance. It materializes this principle by synthesizing two orthogonal signals—a database’s proven, historical efficiency and its potential for future performance gains—allowing SAM to make sound decisions in complex, dynamic environments.
@@
At a macro-level, as shown in Figure 2, AURA employs an Adaptive Active Set (AAS) framework, which operates as a two-state machine. AURA spends the majority of its cycles in the computationally efficient State 1 (Localized Optimization), focusing only on a small subset of tenants ($k\ll K$) to make rapid, high-quality decisions. It transitions to the more comprehensive State 2 (Global Scan) only when a strategic opportunity for improvement is detected (i.e., the active set converges). This design decouples decision latency from the system scale, providing excellent scalability while maintaining decision quality.
@@ 4.2. Algorithm Specification
We now provide the complete specification of the AURA control policy, translating the high-level architecture from the previous section into a concrete implementation. We begin by laying out the necessary data structures and state variables, followed by a detailed walkthrough of the main decision workflow and its innovative components as visually summarized in Figure 2.
@@
The AURA policy maintains state at two levels. The Global Coordinator State ($S_{g}$) includes: (i) Active Set State, which contains the current_active_set of tenants being optimized and cached candidate pools (top_candidates, bottom_candidates); (ii) Convergence State, such as a score_impro-
@@
AURA’s autonomic control policy employs an adaptive two-state architecture for computational efficiency. At the start of each cycle, a key decision is made (Algorithm 1, Line 1): should the system enter the infrequent, comprehensive Global Scan state, or continue in the lightweight, default Localized Optimization state? This decision is governed by a convergence metric that determines when the optimization potential of the current active set has been exhausted. This is achieved by monitoring a sliding window of performance score improvements; if the improvement rate stagnates, convergence is triggered.
If a global scan is triggered (e.g., on a cold start or after convergence), the system performs a highly efficient, linear-time pass using a two-way heap filter to gather fresh candidates (Line 1). After passing an equilibrium check, a knee-point detection algorithm robustly determines the optimal size for a new active set (Line 1). The ComposeSet function(Line 1) then forms the new active set primarily from the TopK candidates, while strategically including some BottomK candidates to identify underperforming tenants whose resources can be proactively reclaimed.
However, the majority of cycles are spent in the fast-path Localized Optimization state. Here, the computationally intensive scoring and allocation adjustments are confined to the small, $O(k)$ active set via the OptimizeInActiveSet function (Line 1). This two-state architecture is the key to AURA’s ability to provide high-quality decisions at a near-constant low cost, regardless of system scale. The core components used within the OptimizeInActiveSet function—our method for scoring tenants by synthesizing their historical efficiency and future potential—are detailed below.
@@ 9    $k_{\text{demand}}\leftarrow\operatorname{FindKneePoint}(TopK)$
Algorithm 1 Algorithm for the AURA Control Policy
@@
The Meta-Adaptive Policy ($\alpha_{t}$). To dynamically balance the $\mathcal{H}$-factor (exploitation) and $\mathcal{V}$-factor (exploration), AURA employs a meta-adaptive policy that computes the weight $\alpha_{t}$. This policy acts as a system-wide “confidence dial”, driven by the insight that the statistical variance of $\mathcal{V}$-factors across the active set serves as a robust proxy for environmental uncertainty. A normalized variance metric, $\kappa$, is calculated; a high $\kappa$ signifies a volatile state with contradictory signals—such as those arising from simultaneous cache warming and workload shifts—pushing the system to be more conservative by increasing the weight on the stable $\mathcal{H}$-factor. Conversely, a low $\kappa$ indicates a stable state with reliable signals, allowing the system to be more opportunistic by trusting the $\mathcal{V}$-factor. The final value is temporally smoothed to ensure the policy itself remains stable.
@@
To understand how the composite scoring function translates into behavior, we analyze its response to four key database archetypes. (i) The Saturated, High-Contribution Instance: This instance maintains a high score due to its large $\mathcal{H}$-factor (from high ops and hit rate). However, its $\mathcal{V}$-factor converges to zero as additional cache yields no further improvement ($\Delta HR\approx 0$), correctly preventing it from wastefully accumulating more resources. (ii) The High-Potential, Low-Resource Instance: An emerging hotspot is characterized by a high marginal gain ($\mathcal{V}$-factor), as small cache additions yield large hit rate improvements. Our normalization mechanism identifies this strong signal, boosting its score and allowing it to rapidly acquire the resources it needs. (iii) The Cache-Inefficient Workload: A “cache polluter” with high ops but persistently low hit rate receives a near-zero score. This is because its $\mathcal{H}$-factor is suppressed by the low hr_norm, and its $\mathcal{V}$-factor is zero, leading to the swift reclamation of its cache. (iv) The Quiescent Instance: An idle database has near-zero ops and no activity, causing both its $\mathcal{H}$-factor and $\mathcal{V}$-factor to be zero. Consequently, it receives no elastic cache, preserving resources for active instances. Besides, this scoring model inherently promotes fairness, as the Global Scan periodically re-evaluates all tenants, giving any instance whose workload changes a chance to enter the active set, while the $\mathcal{H}$-factor ensures that even a low-traffic instance with a high hit rate maintains a non-zero score, preventing it from being permanently starved.